OpenStudy (anonymous):

what is the equation in point slope formof a line that has a slope of 6 and passes through the points (-1-8)

OpenStudy (ryandeutscher):

if i help can you help me

OpenStudy (anonymous):

yes

OpenStudy (ryandeutscher):

whats the question about

OpenStudy (calculusxy):

Do you mean: What is the equation in point slope form of a line that has a slope of 6 and passes through the points \((-1, 8)\) or \((-1, -8)\).

OpenStudy (dayakar):

r u there

OpenStudy (dayakar):

equation of the line whose slope (m) = 6 passing through the point (x1, y1) =(-1,-8) is y-y1= m(x-x1) substitute the values u get the equation, can u do it

OpenStudy (dayakar):

y=(-1) = 6 [x-(-1)] y+1 = 6(x+1) y+1 = 6x+6 0= 6x-y +6-1 0=6x-y +5 therefore 6x-y+5=0

OpenStudy (dayakar):

can u get it